For report 2, do the following:

1. Is the CSTR stable under the following nominal conditions?

Coolant flow  450 kmol/hr
Feed Temperature  24 ^{o}C
Propylene Oxide in Feed 20 kmol/hr

Perform a linear stability analysis to answer the above question.

2. Design a distillation column in CHEMCAD that takes the reactor outlet under nominal conditions and gives a bottoms product that is 99.5% propylene glycol. Show the CHEMCAD simulation results.

3. What happens in the CHEMCAD simulation when the propylene glycol in the feed to the column is increased or decreased by 10%?

4. Suppose the CSTR is operating at a coolant flow rate of 450 kmol/hr and a propylene oxide feed rate of 20 kmol/hr. The initial feed temperature is 20 ^{o}C. Design a PI controller in MATLAB that keeps the outlet proplyne glycol concentration at 0.50 kmol/m^{3}. Show the plots of the feed temperature versus time and the proplyne glycol concentration in the outlet versus time.

5. Implement the above controller in CHEMCAD.